# An improved tabular data generator with VAE-GMM integration

### v1

@@ Abstract
The rising use of machine learning in various fields requires robust methods to create synthetic tabular data. Data should preserve key characteristics while addressing data scarcity challenges. Current approaches based on Generative Adversarial Networks, such as the state-of-the-art CTGAN model, struggle with the complex structures inherent in tabular data. These data often contain both continuous and discrete features with non-Gaussian distributions. Therefore, we propose a novel Variational Autoencoder (VAE)-based model that addresses these limitations. Inspired by the TVAE model, our approach incorporates a Bayesian Gaussian Mixture model (BGM) within the VAE architecture. This avoids the limitations imposed by assuming a strictly Gaussian latent space, allowing for a more accurate representation of the underlying data distribution during data generation. Furthermore, our model offers enhanced flexibility by allowing the use of various differentiable distributions for individual features, making it possible to handle both continuous and discrete data types. We thoroughly validate our model on three real-world datasets with mixed data types, including two medically relevant ones, based on their resemblance and utility. This evaluation demonstrates significant outperformance against CTGAN and TVAE, establishing its potential as a valuable tool for generating synthetic tabular data in various domains, particularly in healthcare.
@@ 1 Introduction
Recent years have witnessed a notable increase in research on synthetic data generation techniques. The past decade has seen the publication of numerous deep learning (DL)-based generative models for homogeneous data types such as images, audio, or text [1, 2, 3, 4, 5, 6]. However, research [7] suggests that the generation of heterogeneous tabular data remains a significant challenge within DL. Tabular datasets are characterized by dense and sparse features, weaker or more complex correlations compared to spatial or semantic data. Additionally, tabular data often include missing values, outliers, and inconsistencies. Furthermore, preprocessing methods used for Deep Neural Networks (DNNs) can lead to information loss. This presents a critical challenge considering the widespread use and importance of tabular data in various fields.
@@
  • While TVAE assumes a Gaussian latent space for data sampling, our model employs a Gaussian Mixture model (GMM) within the VAE architecture. This avoids the potential limitation of a strictly Gaussian distribution imposed by the training process divergence term in VAEs. By leveraging the inherent distribution approximation capabilities of GMMs, we aim to achieve a more accurate representation of the latent space compared to TVAE.
  • Our model outperforms existing solutions by capturing complex feature interactions using diverse distributions that enable handling of continuous and discrete data.
@@ 2.1 Variational Autoencoder
Due to the inherent difficulty of estimating the true posterior density, VAEs employ variational methods. These methods involve defining an approximation, a variational distribution $q_{\phi}(z|x)$ parameterized by $\phi$, and optimizing its parameters to closely resemble the true posterior, as shown in Fig. 1. This optimization typically involves maximizing the Evidence Lower Bound (ELBO), which provides a lower bound on the data’s marginal log-likelihood. For any sample $x_{i}$, the ELBO expression is:
@@
The VAE implementation consists of an encoder and a decoder, both implemented using deep neural networks (DNNs). The encoder approximates the true posterior distribution $p(z|x)$ with a variational distribution $q_{\phi}(z|x)$, parameterized by $\phi$. The parameters of this distribution are obtained by applying a deterministic transformation $g_{\phi}$ to the input data $x$ combined with random noise $\epsilon$. The resulting latent variable $z$ serves as input to the decoder, which aims to reconstruct the original data by estimating the conditional distribution $p_{\theta}(x|z)$, parameterized by $\theta$.
@@ 2.3 Contribution
Previous models such as TVAE make the implicit assumption that the latent space $z$ aligns with the prior Gaussian distribution and therefore to generate new samples of z, sampling from the prior distribution, which is isotropic Gaussian, suffices. The reason is the presence of the KL divergence term in the loss function 1. However, this assumption may not hold because of the presence of complex dependencies and correlations in real-world data. These factors are represented by the other term in the loss function, the marginal log-likelihood distributions. This term can potentially push the latent space away from a Gaussian distribution. In other words, the VAE loss 1 has two components: the log-likelihood of the reconstructed data, which allows the reconstruction of samples from the latent space with high fidelity, and the KL divergence, which acts as a regularizer in the latent space. Note that the KL divergence limits the complexity of the latent space, so it can be thought of as the allowed latent representation complexity. As it is typical to use an isotropic Gaussian as prior, due to the KL being analytical, there is an equilibrium between having a good reconstruction with a low-complexity latent space representation. In practical terms, the actual latent space learned need not be an isotropic Gaussian (as we will show experimentally), and thus sampling $z$ from the prior may not provide the best results in terms of generating new samples.
Therefore, our key contribution lies in employing a BGM for the sampling process of new tabular data. Unlike TVAE, our model does not assume that the actual latent space follows the isotropic Gaussian prior: instead, we additionally model $z$ as a mixture of $K$ Gaussian distributions, each characterized by its mean, covariance matrix, and mixing coefficient. Notably, if $z$ is truly an isotropic Gaussian, the BGM would still be able to effectively approximate it, as it corresponds to the case where $K=1$.
@@ 3 Experimental analysis
Figure 3: Latent space comparison among the three approaches. 300 samples from each latent space were selected to optimize the visualization of the results. The VAE’s latent space $z$ is shown in blue, green is the TVAE’s latent space, and our proposed model’s is in orange. The BGM-modeled latent space closely aligns with $z$ compared to the TVAE’s. This demonstrates the importance of BGM for capturing the latent space distribution and potentially leading to higher-quality generated samples.
@@
  • STD: This dataset deals with Sexually Transmitted Diseases (STD) information from 877 patients. The dataset was downloaded from KMsurv [23] R package. It presents 23 features with various data types (categorical, integer, binary). The target variable again focuses on survival information.
@@
Let us now confirm that intuition with benchmarks on the data. Due to the lack of standardized metrics, the evaluation of data generation models requires a multifaceted approach. We use a combination of resemblance and utility-based evaluations. Resemblance is assessed through a Random Forest (RF) discriminator that measures the similarity of the joint distribution between real and synthetic data. In an ideal scenario of high model performance, the RF accuracy should be close to a random guessing rate (around $0.5$), indicating that it is not able to distinguish between real and synthetic data. Furthermore, we use the technique proposed in [20] to validate the marginal distributions and the correlation between pairs of columns. Utility evaluation, tailored to the type of task, involves the Cox Proportional Hazards (Coxph) model [24] and the Concordance Index (C-index) [25] for SA, and the aforementioned RF discriminator for classification tasks. This evaluation considers two scenarios: training and testing just on real data (benchmark) and training on synthetic data with real data testing. This comprehensive assessment ensures that the generated data exhibit high resemblance (capturing feature distributions and correlations) and utility (supporting robust statistical inference). To benchmark our model’s performance, we compared it with CTGAN and TVAE, the current state-of-the-art models for synthetic data generation. For a coherent comparison, TVAE and our model are implemented in the same way, but we add the BGM sampling to ours. All models were evaluated using the same methodology for a fair and consistent comparison.
@@
Accuracy and 99% confidence interval. Lower is better.
Best values are in bold.

### v2

@@ Abstract
The rising use of machine learning in various fields requires robust methods to create synthetic tabular data that preserve key characteristics while mitigating data scarcity challenges. State-of-the-art approaches, such as CTGAN and TVAE, face difficulties with the intricate structures inherent in tabular data, which often comprise continuous and discrete features with non-Gaussian distributions. To address these limitations, we propose a novel approach based on Variational Autoencoders (VAEs) enhanced with a Bayesian Gaussian Mixture (BGM) model. Unlike other methods that alter the Gaussian prior of the VAE, our approach trains the VAE conventionally and then applies the BGM model to the learned latent space. This allows for a more accurate representation of the underlying data distribution during data generation. Moreover, our model offers enhanced flexibility by accommodating various differentiable distributions for individual features, enabling the handling of continuous and discrete data types. Thorough validation on three real-world datasets, including medically relevant ones, demonstrates significant outperformance compared to CTGAN and TVAE. Our model shows promise as a valuable tool for synthetic tabular data generation across diverse domains, particularly in healthcare.
@@ 1 Introduction
Recent years have witnessed a notable increase in research on synthetic data generation techniques. The past decade has seen the publication of numerous deep learning (DL)-based generative models for homogeneous data types such as images, audio, or text [1, 2, 3]. However, research [4] suggests that the generation of heterogeneous tabular data remains a significant challenge within DL. Tabular datasets are characterized by dense and sparse features, weaker or more complex correlations compared to spatial or semantic data. Additionally, tabular data often include missing values, outliers, and inconsistencies. Furthermore, preprocessing methods used for Deep Neural Networks (DNNs) can lead to information loss. This presents a critical challenge considering the widespread use and importance of tabular data in various fields.
@@
  • While TVAE assumes a Gaussian latent space for data sampling, our model employs a Gaussian Mixture model (GMM) within the VAE architecture. This avoids the potential limitation of a strictly Gaussian distribution imposed by the training process divergence term in VAEs [13].
  • While prior research explored non-Gaussian latent spaces in VAEs [14], [15], these approaches require modifications to the training process, increasing complexity. In contrast, our model leverages the learned latent representation after conventional VAE training and subsequently applies a GMM. Thus, we are able to generate better synthetic data while keeping the VAE training complexity, adding only the BGM after the VAE has been trained
  • We propose a more flexible GMM compared to [16] where a fixed number of components is used in a pre-trained VAE limited to non-tabular data. Our GMM learns the optimal number of components directly from the tabular data.
@@ 2.1 Variational Autoencoder
Due to the inherent difficulty of estimating the true posterior density, VAEs employ variational methods. These methods involve defining an approximation, a variational distribution $q_{\phi}(z|x)$ parameterized by $\phi$, and optimizing its parameters to closely resemble the true posterior (Fig. 1). This optimization typically involves maximizing the Evidence Lower Bound (ELBO), which provides a lower bound on the data’s marginal log-likelihood. For any sample $x_{i}$, the ELBO expression is:
@@
The VAE implementation consists of an encoder and a decoder, both implemented using DNNs. The encoder approximates the true posterior distribution $p(z|x)$ with a variational distribution $q_{\phi}(z|x)$, parameterized by $\phi$. The parameters of this distribution are obtained by applying a deterministic transformation $g_{\phi}$ to the input data $x$ combined with random noise $\epsilon$. The resulting latent variable $z$ is input to the decoder, which aims to reconstruct the original data by estimating the conditional distribution $p_{\theta}(x|z)$, parameterized by $\theta$.
@@ 2.3 Contribution
Previous models such as TVAE implicitly assume that the latent space $z$ aligns with the prior Gaussian distribution. Therefore, to generate new samples of $z$, sampling from the prior distribution, which is isotropic Gaussian, suffices. The reason is the presence of the KL divergence term in the loss function 1. However, this assumption may not hold because of complex dependencies and correlations in real-world data. These factors are represented by the other term in the loss function, the marginal log-likelihood distributions. This term can push the latent space away from a Gaussian distribution. In other words, the VAE loss 1 has two components: the log-likelihood of the reconstructed data, which allows the reconstruction of samples from the latent space with high fidelity, and the KL divergence, which acts as a regularizer in the latent space. The KL divergence limits the complexity of the latent space, so it can be thought of as the allowed latent representation complexity. As it is typical to use an isotropic Gaussian as prior, due to the KL being analytical, there is an equilibrium between having a good reconstruction with a low-complexity latent space representation. In practice, the actual latent space learned need not be an isotropic Gaussian (as we will show experimentally), and thus sampling $z$ from the prior may not provide the best results in generating new samples.
Therefore, our key contribution lies in employing a BGM for the sampling process of new tabular data. Unlike TVAE, our model does not assume that the actual latent space follows the isotropic Gaussian prior: instead, we additionally model the already learned latent space $z$ as a mixture of $K$ Gaussian distributions, each characterized by its mean, covariance matrix, and mixing coefficient. Notably, if $z$ is truly an isotropic Gaussian, the BGM would still be able to effectively approximate it, as it corresponds to the case where $K=1$.
@@ 3 Experimental analysis
Figure 3: Latent space comparison. 300 samples from each VAE, TVAE, and proposed model’s latent space are shown. The BGM-modeled latent space (orange) closely aligns with $z$ (blue) compared to the TVAE’s (green). This demonstrates the importance of BGM for capturing the latent space distribution and potentially leading to higher-quality generated samples.
@@
  • STD: This dataset deals with Sexually Transmitted Diseases (STD) information from 877 patients. The dataset was downloaded from KMsurv [24] R package. It has 23 features with categorical, integer and binary data types. The target variable focuses on survival information.
@@
Let us now confirm that intuition with benchmarks on the data. Due to the lack of standardized metrics, the evaluation of data generation models requires a multifaceted approach. We use a combination of resemblance and utility-based evaluations. Resemblance is assessed through a Random Forest (RF) discriminator that measures the similarity of the joint distribution between real and synthetic data. In an ideal scenario of high model performance, the RF accuracy should be close to a random guessing rate (around $0.5$), indicating that it can not distinguish between real and synthetic data. Furthermore, we use the technique proposed in [21] to validate the marginal distributions and the correlation between pairs of columns. Utility evaluation, tailored to the type of task, involves the Cox Proportional Hazards (Coxph) model [25] and the Concordance Index (C-index) [26] for SA, and the RF discriminator for classification tasks. This evaluation considers two scenarios: training and testing on real data (benchmark) and training on synthetic data with real data testing. This assessment ensures that the generated data exhibit high resemblance (capturing feature distributions and correlations) and utility (supporting robust statistical inference). To benchmark our model’s performance, we compared it with CTGAN and TVAE, the current state-of-the-art models for synthetic data generation. For a coherent comparison, TVAE and our model are implemented in the same way, but we add the BGM sampling to ours. All models were evaluated using the same methodology for a fair and consistent comparison.
@@
Accuracy and 99% CI. Lower is better. Best values are in bold.